OpenStudy (anonymous):

square root of 24x^6

OpenStudy (tkhunny):

This is a tricky one. You can get it wrong. Show us your first attempt.

OpenStudy (igreen):

\(\sf \sqrt{24x^6}\)

OpenStudy (anonymous):

I have no idea how to do this

OpenStudy (igreen):

Okay..first simplify the whole number, do you know how to do that? Simplify \(\sf\sqrt{24}\)

OpenStudy (anonymous):

4.89

OpenStudy (tkhunny):

That is false. You CANNOT have "no idea". That's just wrong. Put away your calculator. This is not a calculator problem.

OpenStudy (anonymous):

I dont know how to do any math without my calculator

OpenStudy (tkhunny):

Simple, then. You cannot solve this problem. You are in the wrong class. You should go have a very thorough discussion with your academic advisor. You're doomed to failure in your present circumstances. Just being honest. Good luck. When you are in the right class, come on back and let's see if we can help you.

OpenStudy (anonymous):

What do you do when you are homeschooled

OpenStudy (anonymous):

because i am

OpenStudy (tkhunny):

Back up and get the stuff you missed. I have told many, many parents over the years: "If you can't handle the math, you must get help or you have no business home-schooling." It's a tough situation. You MUST get local help. You can't do it alone.

OpenStudy (anonymous):

I am homeschooled through K12

OpenStudy (igreen):

To simplify \(\sf \sqrt{24}\), you need to see if any perfect squares are divisible by 24. Do you know what a perfect square is?

OpenStudy (anonymous):

If people are just going to criticize me and not help then they can leave this question and never return

OpenStudy (anonymous):

6

OpenStudy (tkhunny):

No criticism at all. Just the facts. You need more help than can be given , here. We may be able to get you through a problem or two, but you will need MUCH more than that. It is the truth. Seek a solution sooner, rather than later.

OpenStudy (igreen):

No..perfect squares are numbers where, when you square them, you get a whole number. It's basically multiplying whole numbers to themselves. 1 * 1 = 1 Therefore 1 is a perfect square. 2 * 2 = 4 Therefore 4 is a perfect square. 3 * 3 = 9 Therefore 9 is a perfect square. 4 * 4 = 16 Therefore 16 is a perfect square. Now look at these numbers, can you divide any of them to 24 and end up with a whole number? Exclude 1 because that doesn't help us.

OpenStudy (igreen):

when you find the square root of them*

OpenStudy (anonymous):

4

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Correct, so we can separate that from \(\sf \sqrt {24}\) 24 / 4 = ?

OpenStudy (anonymous):

6

OpenStudy (igreen):

Yes.. so we can write: \(\sf \sqrt 4 \times \sqrt 6\) And we simplify the square root of 4: \(\sf 2 \times \sqrt 6\) \(\sf 2\sqrt 6\)

OpenStudy (igreen):

So we have: \(\sf 2 \sqrt{6x^6}\)

OpenStudy (igreen):

Now let's simplify \(\sf \sqrt{x^6}\). By rule: \(\sf \sqrt{a} \rightarrow a^{\frac{1}{2}}\) So \(\sf \sqrt{x^6} \rightarrow (x^6)^{\frac{1}{2}}\)

OpenStudy (igreen):

Now use the exponent rule. If you have an exponent outside of a parenthesis, multiply it to all the exponents inside the parenthesis. What do you get? @Nevaeh_Loves1999

OpenStudy (anonymous):

3

OpenStudy (phi):

If you have not seen Khan's videos, try (for example) https://www.khanacademy.org/math/algebra-basics/core-algebra-foundations/square-roots-for-college/v/understanding-square-roots you can search his site for videos on any math you are studying (or review stuff that you do not understand)

OpenStudy (igreen):

Yes..1/2 * 6 is equal to 3. That gives us: \(\sf x^3\) So our final answer is: \(\sf 2x^3 \sqrt 6\)

OpenStudy (anonymous):

OMG Thank You soooo much

OpenStudy (anonymous):

\[2.2 = 4 \ so \ its \ squareroot \ is\ 2. \\ 3.3 = 9 \ so \ its \ squareroot \ is \ 3 \\ 2.2.2 = 4.2 \ we \ know \ squareroot \ of \ 4 \ but \ what \ is \ for \ 2\] So 2 has decimal square root. So for any given number find all of its factor? factor of 4 is 2.2. factor of 15 is 3.5 How to find factor. Divide the number by 2 first if its divisible then get the quotient after dividing. Then divide that number from 3 so on until the quotient is less than equal to divisor.. eg - 21 , its not divisible by 2, try with 3, its divisible by 3, so now we have 3 as factor and 7 as quotient, divide 7 by 4 failed, by 5 failed, by 6 failed, by 7 success. So quotient is 1 and factor is 3 and 7. As the current number to divide is > quotient stop the procedure. As both 3 and 7 has decimal square root then answer will be decimal or it will not come out of the root.

OpenStudy (igreen):

No problem.